Operations Report – May 15th

Crew 281 Operations Report 15-05-2023
SOL: 1
Name of person filing report: Megan Kane
Non-nominal systems:
· Lower deck shopvac
· Electrical outlets in kitchen
· Remote observatory
· EVA Helmet – One of the helmets has some damage to the collar, base and back. The spare helmet will be used instead
· Headset 4 is non-operational and requires replacement as described EVA report
Notes on non-nominal systems:
· Lower deck shopvac filter is beyond end of life and needs replacement; significant amount of hair and dust by-passed filter and entered blower. This has been mostly removed with pliers. Makeshift filter from previous crew removed from outlet, older filter banged and blown clean, but frame is bent out of shape and cannot easily be reformed. Recommend replacement.
· Electrical outlets in kitchen remain non functional from previous crew.
· Remote observatory still out of commission from previous mission.
· EVA Helmet collar damage similar to previous crew, additional fracture in rear panel, additional use will likely extend fracture.
Spirit rover used: Yes
Hours: (before EVA): 222.2
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 88%
Currently charging: Yes
Opportunity rover used: Yes
Hours: 129.4
Beginning charge: 100%
Ending charge: 99%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 239.8
Beginning charge: 100%
Ending charge: 88%
Currently charging: Yes
Perseverance rover used: Yes
Hours: 264.3
Beginning charge: 100%
Ending charge: 83%
Currently charging: Yes
General notes on rovers:
Summary of Hab operations: put notes here
WATER USE: 65.35 gallons (sols 0 and 1 combined)
Water (static tank): 484.65 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no or yes: no
Summary of internet:
· Password to Outpost Internet has the incorrect capitalization written, resolved by radio communications with Mission Control
· Additional MDRScampus-guest was requested and the access was extended to be from 6-10 pm for additional research resources that cannot be downloaded or synced.
Summary of suits and radios: All suits and radios are fully recharged after todays EVAs
Summary of GreenHab operations: put notes here
WATER USE: 3.43 gallons
Heater: Off
Supplemental light: Off
Harvest: report crop and mass in grams and any other operational notes
Tomatoes 60 harvested for a total of 252 grams and an Average Mass of 0.238095g per tomato.
Installed passive watering system.
Summary of ScienceDome operations: put notes here
Dual split: Heat or AC, On or Off
Summary of RAM operations: (Tools used, work done)
· Water came in around the edges of the skylight window. Everything in the RAM was moist. This is the most likely cause of the rust visible on the tools.
· Removed masking tape for use in the main hab. It is still in the main hab so it remains dry for future use.
· Removed multimeter to lower deck for use in testing comm and suit batteries
· Removed tape measure to lower deck for use in measuring water usage

Summary of any observatory issues: NA
Summary of health and safety issues:
· No injuries,
· hydration packs have been used,
· Anticipate heat and overheating will be the largest challenge going forward
Questions, concerns and requests to Mission Support:
· Downstairs shopvac was clogged and had to be cleaned prior to use. Suggest that future crews empty all vacuums at the end of the mission. Additionally, recommend that filters are checked between missions.

Copyright © The Mars Society. All rights reserved. | Main Site